# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

### v1

@@ 2.2 Frequentist guarantees
The “true” Poisson compound model for $(Y_{n})_{n\geq 1}$ is the model under which, for any $n\geq 1$: i) $Y_{n}$ given $\theta^{\ast}_{n}$ is Poisson distributed with parameter $\theta^{\ast}_{n}$; ii) the $\theta^{\ast}_{n}$’s are independent and identically distributed from the “true” prior distribution, say the oracle prior $G^{\ast}$ with probability mass function $g^{\ast}$ on $\Theta_{d}$. If $G^{\ast}$ was known, then we could estimate $\hat{\theta}_{G}(y)$ by $\hat{\theta}_{g^{*}}(y)$, namely the plug-in estimate obtained from (1) by replacing $g$ with $g^{\ast}$. Under the “true” model, the Bayes risk of $\hat{\theta}_{g^{*}}(y)$ is
$\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})=E^{*}(\hat{\theta}_{g^{*}}(Y_{n})-\theta_{n}^{*})^{2}=\sum_{j=1}^{d}\sum_{y\in\mathbb{N}_{0}}^{\infty}(\hat{\theta}_{g^{*}}(y)-\vartheta_{j})^{2}k(y\mid\vartheta_{j})g^{*}(\vartheta_{j}),$
with $E^{*}$ denoting the expected value with respect to the true model. Not knowing the oracle $g^{*}$, we employ the estimate $\hat{\theta}_{g_{n}}(y)$ in (7). In particular, under the “true” model, the Bayes risk of $\hat{\theta}_{g_{n}}(y)$ is
$\mathcal{R}(g^{*},\hat{\theta}_{g_{n}})=E^{*}(\hat{\theta}_{g_{n}}(Y_{n})-\theta^{*}_{n})^{2}=\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})+E^{*}(\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g^{*}}(Y_{n}))^{2}.$
The Bayesian regret (of not knowing $g^{*}$ and using $g_{n}$ instead) is the difference between $\mathcal{R}(g^{*},\hat{\theta}_{g_{n}})$ and $\mathcal{R}(g^{*},\hat{\theta}_{g^{*}})$, i.e.
$\text{Regret}(g_{n},g^{*})=E^{*}((\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g^{*}}(Y_{n}))^{2}=\sum_{j=1}^{d}\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g^{*}}(y))^{2}k(y\mid\vartheta_{j})g^{*}(\vartheta_{j}).$ (9)
See Efron (2019) for details on (9). The next theorem provides the large $n$ asymptotic behaviour of the regret.
@@
Let $g_{n}$ be as in (5) with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1)$, and let $g^{\ast}$ be the probability mass function of the oracle prior $G^{\ast}$. Denote by $\mathds{P}^{*}$ a probability measure under which the “true” Poisson compound model for $(Y_{n})_{n\geq 1}$ holds true. Then, for every $\delta<2-1/\gamma$, it holds that
$\text{Regret}(g_{n},g^{*})=o(n^{-\delta})\quad\mathds{P}^{\ast}\,\text{-}\,a.s.$
See Supplementary Material C.3 for the proof of Theorem 2.4. The regret analysis of Theorem 2.4 is provided under the standard choice $\alpha_{n}=(1/(\alpha+n))^{\gamma}$, though other learning rates may be considered.
@@ 3 Numerical illustrations
We validate empirically our sequential $g$-modeling on synthetic and real data. We compare the quasi-Bayes (NP-QB) estimate $\hat{\theta}_{g_{n}}$ of $\hat{\theta}_{G}$, at fixed $n\geq 1$, with the estimates obtained by the following methods: i) nonparametric empirical Bayes (NP-EB) (Robbins, 1956); ii) nonparametric maximum likelihood (NP-ML) (Jiang and Zhang, 2009); iii) nonparametric minimum squared Hellinger distance (NP-MD) (Jana et al., 2024); parametric Bayes empirical Bayes (P-EB) with a Gamma hyperprior on $G$ (Deely and Lindley, 1981). NP-EB is the sole to produce a closed-form estimate, i.e. the estimate obtained by replacing $p_{G}$ in (1) with the empirical distribution (2). NP-ML, N-PMD and P-EB require exact or approximate optimization algorithms: i) NP-ML and N-PMD estimates are obtained from Jana et al. (2024, Algorithm 1); ii) the P-EB estimate is obtained from Deely and Lindley (1981, Equation 4.4), with the parameters of the Gamma hyperprior estimated by maximum (marginal) likelihood. The empirical performance of the above estimates is evaluated by the root mean squared error (RMSE) and the mean absolute deviation (MAD).
@@ 3.1 Synthetic data
We set $n=50,100,200,500$, and generate $(Y_{1},\ldots,Y_{n})$ under the Poisson compound model, with $(\theta_{1},\ldots,\theta_{n})$ i.i.d from a Gamma prior $G$ with scale $2$ and shape $4$. For NP-QB, we set Newton’s algorithm as follows: i) $\theta$ takes value in $\Theta=\{0.01,0.02,\ldots,99.99,100\}$; ii) the number $d$ of mixture components is the cardinality of $\Theta$, i.e. $d=10,000$; iii) the initial guess $g_{0}$ is the Uniform distribution on $\Theta$; iv) the learning rate is $\alpha_{n}=(1+n)^{-1}$. Results are displayed in Figure 1 and Table 1.
Figure 1: Data $Y_{i}$’s against the true $\theta_{i}$’s (black), and estimates of the $\theta_{i}$’s: NP-EB (green), NP-ML (cyan), NP-MD (magenta), P-EB (red) and NP-QB (blue).
Table 1: Values of RMSE and MAD for the NP-EB, NP-ML, NP-MD, P-EB and NP-QB estimates.
NP-EB NP-ML NP-MD P-EB NP-QB
RMSE   $n=100$ 6.184 2.230 2.253 2.110 2.235
RMSE   $n=500$ 3.737 2.238 2.228 2.244 2.236
MAD     $n=100$ 4.100 1.722 1.731 1.632 1.719
MAD     $n=500$ 2.459 1.773 1.768 1.782 1.774
@@ 3.2 Real data
We consider a dataset of tweets with at least 50 retweets on Twitter in the period October 7 to November 6, 2011 (30 days). The dataset is available at https://snap.stanford.edu/seismic/, and it consists of $166,783$ tweets that generate $34,617,704$ retweets. The data provides information on the tweet ID, the tweet post time, the retweet post time and the number of followers of the poster/retweeter. For each tweet, we focus on the number of retweets generated within $30$ seconds from the tweet post time; in principle, the faster a retweet is, the more likely it was generated by a chatbot artificial intelligence rather than a human. The count data of fast retweets arrive at high-velocity, thus motivating the need of an online fast processing. From the first two rows of Table 2, $40,259$ tweets have 00 retweets in $30$ seconds, $28,339$ tweets have $1$ retweet in $30$ seconds, etc. Interest is in estimating the number of retweets expected for a tweet. Table 2 displays the estimates obtained through the application of NP-EB, NP-ML, NP-MD, P-FB and NP-QB.
Table 2: Counts $n_{y}$ of the number of tweets that have $y$ retweets within $30$ seconds from the tweet post time, for $y=0,1,\ldots,7$, and corresponding estimates of the number of retweets expected.
@@
NP-QB 0.607 0.818 1.632 3.076 4.007 4.519 5.023 5.707
We refer to Supplementary Material D.2 for a version of Table 2 with more values of $y$, which shows that NP-QB competes with NP-ML, NP-MD and P-EB, and all these methods outperform NP-EB. The Supplementary Material D.2 includes another application to same Twitter dataset, but with respect to the number of followers generated by fast retweets. See also Supplementary Material D.3 for an application to the benchmark auto accident dataset (Efron and Hastie, 2021, Table 6.1).
@@ 4 Discussion
The Supplementary Material E contains an extension of our sequential $g$-modeling to the multidimensional Poisson compound decision problem (Brown et al., 1985; Johnstone, 1986). Further extensions may be considered with respect to compound decision problems beyond the Poisson setting, with applications to high-dimensional problem, e.g. exponential families (Zhang, 2003). Sequential $f$-modeling is also an interesting direction for future research (Brown et al., 2013).
@@ C Proofs for Section 2
C.1 Proof of Theorem 2.1
@@
It follows by Theorem B.2 that $M_{n,\infty}=b_{n}^{1/2}(\langle g_{n}-\tilde{g},\mathbf{u}\rangle)$ converges to $\mathcal{N}(0,\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u})$ in the sense of almost-sure conditional convergence. Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(g_{n}-\tilde{g})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,V_{\tilde{g}})$.
$\sum_{n=1}^{\infty}\alpha_{n}^{4}b_{n}^{2}=\sum_{n=1}^{\infty}\left(\frac{1}{\alpha+n}\right)^{4\gamma}(2\gamma-1)^{2}\alpha^{2\gamma}n^{4\gamma-2}<+\infty,$
$b_{n}\sup_{k\geq n}\alpha_{k}^{2}=(2\gamma-1)n^{2\gamma-1}\left(\frac{1}{\alpha+n}\right)^{2\gamma}\rightarrow 0,$
and
$b_{n}\sum_{k\geq n}{\alpha_{k}^{2}}=(2\gamma-1)n^{2\gamma-1}\sum_{k\geq n}\left(\frac{1}{\alpha+k}\right)^{2\gamma}\rightarrow 1.$
@@
The proof is based on Theorem C.2 and Theorem B.5. Let us write
@@
C.3 Proof of Theorem 2.4
The proof of Theorem 2.4 is based on the following lemma C.3 and on the rate of convergence of $g_{n}$ to $g$, under i.i.d. observations. This last result is known in the literature (Martin and Tokdar, 2009), but we think that the proof is not complete. We provide for completeness the proof in Lemma C.5.
@@
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
where $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are the representations of $g_{n}$ and $g^{*}$ in $\mathbb{R}^{d}$.

### v2

@@ 2.3 Frequentist guarantees
The “true” Poisson compound model assumes that the observations $Y_{n}$’s are i.i.d. from a Poisson mixture model with a “true” mixing distribution, say the oracle prior $G^{\ast}$ on $\Theta\subseteq\mathbb{R}^{+}$. If $G^{\ast}$ was known, then we could estimate $\hat{\theta}_{G}(y)$ in (1) with the oracle $\hat{\theta}_{G^{*}}(y)$. Instead, not knowing the oracle prior $G^{*}$, we employ the estimate $\hat{\theta}_{g_{n}}(y)$ in (5). Here, Proposition 2.1 is leveraged in order to establish asymptotic frequentist guarantees of $\hat{\theta}_{g_{n}}(y)$ under the “true” model for $(Y_{n})_{n\geq 1}$, demonstrating consistency and asymptotic optimality in terms of vanishing excess Bayes risk or regret.
Let $G^{\ast}$ be an oracle prior on $\Theta\subseteq\mathbb{R}^{+}$ that satisfies the assumptions of Proposition 2.1, and, for a fixed $\eta>0$, let $g^{\ast}_{\eta}$ be its $d_{\eta}$-KL-discretization. We define the regret of using $\hat{\theta}_{g_{n}}(y)$ instead of $\hat{\theta}_{g_{\eta}^{*}}(y)$ as
$\text{Regret}(g_{n},g_{\eta}^{*})=E_{G^{*}}((\hat{\theta}_{g_{n}}(Y_{n})-\hat{\theta}_{g_{\eta}^{*}}(Y_{n}))^{2}=\int_{\Theta}\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g_{\eta}^{*}}(y))^{2}k(y\mid\theta)G^{\ast}(\mathrm{d}\theta)$ (7)
with $E_{G^{*}}$ being the expected value with respect to the “true” model; see Efron (2019, Section 3) for details on (7). We denote by $\mathds{P}_{G^{*}}$ the probability measure under which the “true” model for $(Y_{n})_{n\geq 1}$ holds true.
@@
Let $G^{\ast}$ be an oracle prior on $\Theta\subseteq\mathbb{R}^{+}$ such that $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G^{\ast}}(y)\leq m_{k}$ for some $k\in\mathbb{N}$ with $k\geq 2$ and $m_{k}\in\mathbb{R}$, and for a fixed $\eta>0$ let $g^{\ast}_{\eta}$ be the $d_{\eta}$-KL-discretization of $G^{\ast}$ on $\Theta_{\eta}$. For each $n\geq 1$, let $g_{n}(\vartheta_{j})$ be as in (3) for $\vartheta_{j}\in\Theta_{\eta}$, with $g_{0}(\vartheta_{j})>0$ for every $j=1,\dots,d_{\eta}$, and with $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$. If $\sum_{y\in\mathbb{N}_{0}}d_{\eta}^{2y}p_{G^{\ast}}(y)<+\infty$ then, for every $y\in\mathbb{N}_{0}$,
$\hat{\theta}_{g_{n}}(y)\rightarrow\hat{\theta}_{g^{\ast}_{\eta}}(y)\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$ (8)
Furthermore, if $\gamma<1$ then for every $\delta<2-1/\gamma$
$\text{Regret}(g_{n},g_{\eta}^{\ast})=o(n^{-\delta})\quad\mathds{P}_{G^{*}}\,\text{-}\,a.s.$ (9)
See C.4 for the proof of Theorem 2.4. Under the “true” model for the observations $Y_{n}$’s, with the oracle prior $G^{\ast}$ belonging to the class of light-tail distributions, Theorem 2.4 establishes that $\hat{\theta}_{g_{n}}(y)$ is a consistent (8) and asymptotically optimal (9) estimate of $\hat{\theta}_{g^{\ast}_{\eta}}(y)$. Although Theorem 2.4 is stated with respect to the standard choice $\alpha_{n}=(\alpha+n)^{-\gamma}$, alternative learning rates may also be considered.
@@ 3 Numerical illustrations
We validate our methodology empirically, by comparing the quasi-Bayes empirical Bayes (QB-EB) estimate $\hat{\theta}_{g_{n}}$, at fixed $n\geq 1$, with estimates from: i) nonparametric empirical Bayes (NP-EB) (Robbins, 1956); ii) nonparametric maximum likelihood (NP-ML) (Jiang and Zhang, 2009); iii) nonparametric minimum squared Hellinger distance (NP-MD) (Jana et al., 2024); parametric Bayes empirical Bayes (P-EB) with a Gamma hyperprior on $G$ (Deely and Lindley, 1981). NP-EB uniquely yields a closed-form estimate by replacing $p_{G}$ in (1) with the empirical distribution, while NP-ML, NP-MD and P-EB require exact or approximate optimization algorithms. NP-ML and NP-MD estimates follow the vertex direction algorithm Jana et al. (2024, Algorithm 1), which, as QB-EB, relies on a discretization of $G$ over a finite grid. P-EB estimates use Deely and Lindley (1981, Equation 4.4), with hyperparameters of $G$ estimated by maximum (marginal) likelihood. Performance is assessed via root mean squared error (RMSE) and mean absolute deviation (MAD).
@@ 3.1 Synthetic data
[…]
Figure 1: Data $Y_{i}$’s against the $\theta_{i}$’s (black circle), and estimates of the $\theta_{i}$’s: NP-EB (green), NP-ML (cyan), NP-MD (magenta), P-EB (red) and QB-EB (blue).
Table 1: Values of RMSE and MAD for NP-EB, NP-ML, NP-MD, P-EB and QB-EB estimates.
NP-EB NP-ML NP-MD P-EB QB-EB
RMSE   $n=100$ 2.516 1.695 1.648 1.478 1.675
RMSE   $n=500$ 2.108 1.341 1.515 1.353 1.348
MAD     $n=100$ 1.807 1.334 1.3 1.179 1.335
MAD     $n=500$ 1.4 1.068 1.197 1.073 1.071
@@ 3.2 Real data
We analyze a dataset of tweets that received at least $50$ retweets on Twitter between October 7 and November 6, 2011, available at https://snap.stanford.edu/seismic/. It contains $166,783$ tweets generating $34,617,704$ retweets, along with metadata such as tweet ID, posting time, retweet time, and follower counts. For each tweet, we focus on the number of retweets received within the first $30$ seconds after posting: fast retweeting is often indicative of automated or bot-driven activity, as opposed to human behavior. The underlying retweet intensity can thus be viewed as a latent measure of a tweet’s instantaneous virality, or more generally its propensity to trigger immediate attention. Identifying tweets with unusually high intensity in real time could help detect automated activity or early-stage virality. Since these counts arrive rapidly in a streaming fashion, they motivate the need for a sequential, online estimation method such as the proposed QB-EB.
Table 2 summarizes the empirical distribution of fast retweet counts, e.g., 40,259 tweets received none, 28,339 received one, etc. We apply NP-EB, NP-ML, NP-MD, P-FB, and QB-EB to estimate the intensity for each tweet, quantifying its fast-retweet potential. Results are shown in Table 2.
Table 2: Counts $n_{y}$ of the number of tweets with $y$ fast retweets, and corresponding estimates.
@@
QB-EB 0.607 0.818 1.632 3.076 4.007 4.519 5.023 5.707
We refer to D.2 for another application of the Twitter dataset, with respect to the followers generated by retweets, and for an application to the benchmark accident dataset (Efron and Hastie, 2021, Table 6.1).
@@ 4 Discussion
An extension of the sequential $g$-modeling methodology to the multidimensional Poisson compound decision problem (Brown et al., 1985) is presented in Section E. While our focus is on the Poisson setting, the methodology is readily extendable to other discrete linear exponential families (Banerjee et al., 2021), as well as to the Gaussian setting (Jiang and Zhang, 2009; Soloff et al., 2024).
@@ C Proofs for Section 2
C.1 Proof of Proposition 2.1
We first prove that there exists a probability mass function $f\in\mathcal{S}_{\eta}$ such that $KL(p_{G}||p_{f})<2\eta.$ Since the $k$-moment of the Poisson distribution with parameter $\lambda$ is greater than $\lambda^{k}$, then the assumption $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G}(y)\leq m_{k}$ implies that
$\int_{0}^{\infty}\theta^{k}G(\mathrm{d}\theta)\leq m_{k}.$
For every $i=1,\dots,d_{\eta}$, let $w_{i}=G((\vartheta_{i-1},\vartheta_{i}])$ with $\vartheta_{0}=0$. Let $F$ be the probability measure
$F=\sum_{i=1}^{d_{\eta}}w_{i}\delta_{\vartheta_{i}}+G(\vartheta_{d_{\eta}},+\infty)\delta_{\vartheta_{d_{\eta}}},$
and let $f$ denote its density with respect to the counting measure on $\Theta_{\eta}$. Then
$F=\sum_{i=1}^{d_{\eta}}\delta_{\vartheta_{i}}\int_{I_{i}}G(\mathrm{d}\theta)+\delta_{\vartheta_{d_{\eta}}}\int_{I_{d_{\eta}+1}}G(\mathrm{d}\theta),$
where $I_{i}=(\vartheta_{i-1},\vartheta_{i}]$ for $i=1,\dots,d_{\eta}$ and $I_{d_{\eta}+1}=(\vartheta_{d_{\eta}},+\infty)$. Since the Kullback-Leibler divergence is convex, then, denoting $M=d_{\eta}\eta$, we can write that
$KL(p_{G}||p_{f})$ $\leq\sum_{y\in\mathbb{N}_{0}}\sum_{i=1}^{d_{\eta}}\int_{I_{i}}\log\left(\frac{e^{-\theta}\theta^{y}}{e^{-\vartheta_{i}}\vartheta_{i}^{y}}\right)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)+\sum_{y\in\mathbb{N}_{0}}\int_{M}^{\infty}\log\left(\frac{e^{-\theta}\theta^{y}}{e^{-M}M^{y}}\right)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)$
$<\int_{0}^{M}\eta G(\mathrm{d}\theta)+\int_{M}^{\infty}\sum_{y\in\mathbb{N}_{0}}y\log(\theta)\frac{e^{-\theta}\theta^{y}}{y!}G(\mathrm{d}\theta)$
$\leq\eta+\int_{M}^{\infty}\theta\log(\theta)G(\mathrm{d}\theta)$
$\leq\eta+\frac{\log M}{M^{k-1}}\int_{M}^{\infty}\theta^{k}G(\mathrm{d}\theta)$
$\leq 2\eta.$
We now prove that the minimization problem $\mbox{\sl argmin}_{f\in\mathcal{S}_{\eta}}KL(p_{G}||p_{f})$ has a unique solution. Representing each probability mass function $f$ with support $\Theta_{\eta}$ as a vector $\mathbf{f}$ taking values in the simplex $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\},$ we can see $\ell(\mathbf{f})=KL(p_{G^{*}}||p_{f})$ as a strictly convex function defined on a convex and compact set. Thus the global minimum $\mathbf{g}_{\eta}$ exists and is unique. $\square$
C.2 Proof of Theorem 2.2
@@
It follows by Theorem B.2 that $M_{n,\infty}=b_{n}^{1/2}(\langle g_{n}-\tilde{g},\mathbf{u}\rangle)$ converges to $\mathcal{N}(0,\mathbf{u}^{T}V_{\tilde{g}}\mathbf{u})$ in the sense of almost-sure conditional convergence. Since this is true for every $\mathbf{u}$ with $||\mathbf{u}||=1$, then $b_{n}^{1/2}(g_{n}-\tilde{g})$ converges in the sense of almost sure conditional convergence to $\mathcal{N}(0,V_{\tilde{g}})$. ∎
@@
The proof is based on Theorem C.2 and Theorem B.5. We first prove that (C.5) holds under the assumptions of Theorem 2.3. We can write that
$b_{n}\uparrow+\infty,\quad\sum_{n\geq 1}\alpha_{n}^{4}b_{n}^{2}<+\infty,\quad b_{n}\sum_{k\geq n}\alpha_{k}^{2}\rightarrow 1.$
In particular, $\alpha_{n}^{4}b_{n}^{2}\rightarrow 0$, as $n\rightarrow\infty$. Since $(\alpha_{n})$ is non increasing, then
$b_{n}\sup_{k\geq n}\alpha_{k}^{2}=b_{n}\alpha_{n}^{2}\rightarrow 0.$
Thus, (C.5) holds. Let us write
@@
C.4 Proof of Theorem 2.4
The proof of Theorem 2.4 is based on lemmas C.3, C.7 and on the rate of convergence of $g_{n}$ to $g_{\eta}^{*}$, under i.i.d. observations. This last result is known in the literature (Martin and Tokdar, 2009), but we think that the proof is not complete. We provide for completeness the proof in Lemma C.5.
In this section, we will occasionally represent probability mass functions $g$ on $\Theta_{\eta}$ as vectors $\mathbf{g}$, taking values either in the simplex $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0\;(i=1,\dots,d_{\eta}-1),\;\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$ or in $\mathbb{R}^{d_{\eta}}$, depending on the context, which will be specified when needed. Similarly, the Poisson kernel is represented as a vector in $\Delta_{\eta}$ or $\mathbb{R}^{d_{\eta}}$, by defining $\mathbf{k}_{y}(\vartheta_{j})=k(y\mid\vartheta_{j})$. We denote component-wise multiplication of vectors by $\circ$.
@@
If
$\sum_{y\in\mathbb{N}_{0}}d_{\eta}^{2y}p_{G^{*}}(y)<+\infty,$ (C.6)
then the $d_{\eta}$-KL-discretization $g_{\eta}^{*}$ of $G^{*}$ satisfies for every $i=1,\dots,d_{\eta}$ the equation in $g$:
$g(\vartheta_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)=0.$ (C.7)